OpenStudy (undeaddragon26):

My Flower Basket delivers flowers within a 25 mile radius of their store. If this area is represented graphically, with My Flower Basket located at (4, 3), what is the equation that represents the delivery area? @Venny

OpenStudy (undeaddragon26):

(x-4)^2 + (y-3)^2 = 625 Center (4,3) Radius = 25 like this

OpenStudy (undeaddragon26):

?

OpenStudy (anonymous):

This is just like the other ones. the equation of a circle is (x−h)^2+(y−k)^2=r^2 here h=4, k=3, and r=25 (x-4)^2+(y-3)^2=25^2

OpenStudy (anonymous):

yes you are exactly right

OpenStudy (anonymous):

sorry 4 breakin in tha converstion but did u read wut I said in history chat undead

OpenStudy (undeaddragon26):

no i can't chat

OpenStudy (anonymous):

so then just pm me n i'll respond here